OpenStudy (anonymous):

How come cos^2u+sin^2u =2cos^2u - 1 = 1- 2sin^2u

OpenStudy (anonymous):

Check the link...will help u http://mathscene.co.uk/t/F/Su57k02.htm

OpenStudy (asnaseer):

should your question be:\[\cos^2(u)-\sin^2(u)=2\cos^2(u)-1=1-2\sin^2(u)\]

OpenStudy (anonymous):

asnaseer I see the Equation button when writing a reply but not an original post...

OpenStudy (asnaseer):

I am not sure what you mean - what are you trying to do?

OpenStudy (anonymous):

Next time I post a question, I'd like the nicer format like yours, however I don't see the equation button when making a new post. I guess this is a technical support question :P

OpenStudy (asnaseer):

Yes - I see what you were asking now. Even though there is no Equation button, if you know how to enter the \(\LaTeX\) equations then you can still do so while typing up the question. Inline equation elements like this \(x^2=1\) are entered by enclosing them between \\((\) and \\()\) characters. Equations entered between \\([\) and \\(]\) will appear on the next line. e.g. \[x^2=1\]is on the next line here.